# An Insight into Security Code Review with LLMs: Capabilities, Obstacles and Influential Factors

### v3

@@ 3.2. Data Collection
3.3. Research Procedure of RQ1
Table 1. Versions and hyper-parameters of 6 LLMs studied.
Model Version Context Length Parameter
@@ 3.3. Research Procedure of RQ1
3.3.1. LLM Selection
To fully evaluate the performance of LLMs in security defect detection, we selected four representative LLMs from the proprietary GPT and Gemini families, along with two from the open-source Llama 2 family, which are the most popularly used LLMs when we started our experiments in December 2023. Specifically, for proprietary LLMs, we chose GPT-3.5 and GPT-4 provided by the ChatGPT platform (May 2024 version), and GPT-4 Turbo and Gemini Pro accessed through the API. For open-source LLMs, we deployed the smallest and largest chat versions of models in Llama 2 series–namely, Llama 2 7B and Llama 2 70B–on a server with eight NVIDIA GeForce RTX 4090 GPUs. Throughout the experiments, we kept all hyper-parameters of these LLMs at their default settings. The versions, token constraints, and parameter settings of LLMs used in our study are summarized in Table 1.
@@ 3.3.1. LLM Selection
3.3.2. Prompt Design
Prompt engineering is a process of designing prompts to optimize model performance on downstream tasks (Liu et al., 2023). To design prompts, we followed the best practices published by OpenAI (OpenAI, 2023) and the three prompt design strategies in Zhang et al. (Zhang et al., 2023): basic prompt, enhanced prompt with auxiliary information, and CoT prompt. During our preliminary exploration stage, we tried prompts with various auxiliary information and different CoT intermediate reasoning steps. To fully harness the capabilities of LLMs in detecting security defects, we finally selected five best-performing prompts. Fig. 2 illustrates the construction strategies for the five prompts.
  1. (1)
@@ 3.3.2. Prompt Design
  4. (4)
    Prompt 4 ($\textbf{P}_{cid}$): $\textbf{P}_{b}$ + Specific CWE Instruction. $\textbf{P}_{cid}$ provides a specific list of the first level CWEs in VIEW-1000 (Research Concepts) [^1] VIEW-1000 is a weaknesses classification framework designed for academic research, widely adopted in prior work (e.g.,  (Pan et al., 2023)). It organizes CWEs into a hierarchical structure based on parent-child relationships among CWE entries. The first level of CWEs in VIEW-1000 encompasses a broad range of commonly encountered security defects, thus applied in Prompt 4 ($\textbf{P}_{cid}$)..
  5. (5)
@@
Figure 2. Construction templates for the five prompts
Table 2. Adjusted IHMU-category for LLM responses.
Classification Definition
Instrumental (I) The response explicitly indicates the existence of the security defect identified by the reviewer and provides a fully accurate description.
Helpful (H) The response raises concerns related to the security defect identified by the reviewer, but their descriptions may not be entirely accurate or specific enough.
Misleading (M) The response does not provide any helpful information, e.g., claiming no defects were found or failing to perform the detection task.
@@
Uncertain (U) The response does not mention any defects identified by the reviewer but instead points out the existence of other security defects. Due to the lack of expertise of the researchers about the code and its context, the actual existence of these security defects cannot be confirmed.
3.3.3. Evaluation Metrics
@@
During the experiments, the responses of LLMs were provided in the form of natural language text, making it necessary to manually inspect these responses to ensure the accuracy of the performance measurements. We considered that LLM-generated responses could be evaluated in a manner similar to the approach adopted by Mahajan et al. (Mahajan et al., 2020) for measuring responses in Stack Overflow. We adapted the evaluation metrics formulated in their work (Mahajan et al., 2020) as follows: First, we excluded the ‘Unavailable’ category defined by Mahajan et al. because LLMs do not have a case of no response. Next, the definitions of each category were adjusted to match the task of security code review. Specifically, each response was categorized into one of the four rating categories detailed in Table 2. The corresponding evaluation metrics are I-Score ($\frac{I}{I+H+U+M}\times 100\%$), IH-Score ($\frac{I+H}{I+H+U+M}\times 100\%$), and M-Score ($\frac{M}{I+H+U+M}\times 100\%$). Higher I-Score and IH-Score indicate a stronger ability of the model to detect security defects identified by reviewers. At the same time, a lower M-Score signifies less misleading results produced during security defect detection. Given that LLMs exhibit considerable randomness, we repeated the experiment for each LLM-prompt combination three times to mitigate the impact of such randomness. We calculated the average of the three performance scores as the final result.
3.3.4. Baselines
We tried different tools and learning techniques for detecting security defects to select appropriate baselines for our work.
@@ 3.3.4. Baselines
To compare with LLMs, we focused on techniques providing fine-grained detection, including classification, localization, and description of defects. This led us to exclude techniques with insufficient detection granularity, such as ProRLearn (Ren et al., 2024), GRACE (Lu et al., 2024) and LineVul (Fu and Tantithamthavorn, 2022). Given that we evaluated the performance of LLMs on the C/C++ and Python datasets separately, for language-specific static analysis tools, we selected CppCheck (Cppcheck Solutions AB, 2024) for C/C++ and Bandit (ban, 2024) for Python. Furthermore, we also used CodeQL (GitHub, 2021) and SonarQube (SonarSource SA, 2024) to analyze Python files. These two tools can directly analyze code files written in dynamically-typed languages like Python but require code files in statically-typed languages like C/C++ to be compiled first. Therefore, we could not apply them to our C/C++ dataset, which is comprised of uncompiled and isolated C/C++ code files. Additionally, since Semgrep (sem, 2024) supports multiple programming languages and can analyze individual code files, we adopted it for both C/C++ and Python datasets. To obtain comprehensive analysis results, CodeQL adopted the Python-security-and-quality.qls test suite, covering a wide range of issues from basic code structure and naming conventions to advanced security and performance vulnerabilities. Other tools utilized their default rule sets for analysis.
3.3.5. Data Labelling
@@
Prompts were constructed using five templates and then fed into six LLMs to collect their responses. Since the responses of LLMs were provided in the form of natural language text, to rate the responses accurately, we manually inspected the content of LLM-generated responses, their corresponding source code, and security defects identified by reviewers to categorize them into the four categories formulated in Table 2. To mitigate bias, pilot data labelling was conducted, in which we randomly selected 100 out of 534 code files from our dataset, constructed five prompts for each file and fed these prompts into LLMs to collect responses. Using the definitions of categories in Table 2 as the labelling criteria, the first and third authors labelled the collected responses separately. Discrepancies in labelling results were discussed with the second author to reach a consensus. The inter-rater reliability was calculated using Cohen’s Kappa coefficient (Cohen, 1960), yielding a value of 0.89. Then, the first and third authors labelled all the rest of the responses and discussed any ambiguous cases with the second author to finalize the categorization. Based on the labelling results, we calculated the performance scores of each LLM-prompt combination in the Python and C/C++ dataset. We found that the combination of GPT-4 and prompt $\textbf{P}_{cid}$ outperformed the others (as detailed in Section 4.1). Therefore, we aggregated the responses generated by GPT-4 with $\textbf{P}_{cid}$ across the Python and C/C++ datasets, resulting in a total of 421 responses (some code files exceeded the token limitation of GPT-4) for each of the three repetitive experiments, to further analyze RQ2 and RQ3.
3.3.6. Consistency Calculation
@@
To measure the randomness of LLMs, we adopted the metrics in Chang et al.’s work (Chang et al., 2024) and utilized entropy to measure the consistency of responses generated by LLMs across three repetitive experiments. Specifically, an LLM generated three responses for a given code file under a specific prompt template. We then calculated the entropy for the four response categories in the three responses. We determined the overall consistency by averaging the entropy across all code files using the following formula: $\frac{1}{R}\sum_{r\in R}\text{entropy}(p^{r}_{1},\dots,p^{r}_{4})$, where R is the response count in each round of the experiment. A higher average entropy suggests the LLM has higher inconsistency in repetitions of experiments.
3.4. Research Procedure of RQ2
@@
We manually inspected the responses of GPT-4 with $\textbf{P}_{cid}$ across the three repetitive experiments to analyze the quality problems present in these responses. We employed the open coding and constant comparative method (Glaser, 1965) with a predefined category by the MAXQDA tool for qualitatively analyzing the responses.
@@
Based on the problem categories formulated by Kabir et al. (Kabir et al., 2024) and the problems of LLM-generated text in various domains identified by previous studies (Borji, 2023; Mitrović et al., 2023; Kalla and Kuraku, 2023), two of the authors firstly inspected a few GPT-4 responses independently and recorded their observations. Then they collaboratively reviewed the inspected responses and discussed to establish an initial codebook, resulting in four themes: Correctness, Understandability, Conciseness, and Compliance, each with a series of problem types. In particular, we excluded the Consistency theme defined in Kabir et al. (Kabir et al., 2024) from our study. The reason is that the consistency between the LLM’s responses and the reviewers’ assessments (i.e., the ground truth) has been investigated in RQ1. Following the codebook, we conducted a pilot data analysis across two iterations to improve the agreement between the two authors. In each round, ten responses were randomly selected for the two authors to analyze their quality problems. The results were then compared and discussed, with partial adjustments made to the codebook. Since a single response could contain multiple problems, two authors completed the remaining data analysis task where the labels are not mutually exclusive. Therefore, we calculated the agreement level using Krippendorff’s alpha (Krippendorff, 2004) with Jaccard distance, which was improved from 0.40 in the first round to 0.80 in the second, indicating a high level of agreement. Then the remaining data analysis was completed by the two authors and verified by the second author. Throughout the entire procedure, we employed a negotiated agreement approach (Campbell et al., 2013): any conflicts were consulted and addressed by three authors, ensuring the reliability of data analysis. For each type of problem, we computed its average frequency in three repetitive experiments as the final occurrence rate of this problem in GPT-4 responses.
3.5. Research Procedure of RQ3
@@
To answer RQ3, we adopted the approach formulated by Harrell (Harrell, 2001) to construct a regression model. From Table 2, we can see that the response variables in our dataset are the ratings of I, H, U, and M. In terms of the helpfulness of the response, by definition, I is more important than H, and so forth. Accounting for the ordinal nature of the response variable and the randomness of the LLM in three rounds of experiments, we ultimately aggregated the data from all three experiments to fit a cumulative link mixed model (Powers and Xie, 2008). This model supports ordered categorical response variables and allows us to include random effects to group responses in each round together. The specific construction process is described in the following subsections.
@@
Table 3. Factors that may influence the performance of LLM in security code review
@@
Token Adjusted from  (Paul et al., 2021; McIntosh et al., 2016) Number of tokens of the code file. Computed using the TikToken (tik, 2024) tokenizer. Larger code files may be more difficult for LLMs to analyze.
@@
SecurityDefectType Transformed from  (Chen et al., 2023) Type of the security defect that the reviewers identified in this code file, i.e., one of the 15 types of security defect formulated by Yu et al. (Yu et al., 2023). ChatGPT is proven to display distinct detection performances across various vulnerabilities (Chen et al., 2023; Zhou et al., 2023). Thus, we assume that LLMs may be more adept at detecting certain specific types of security defects in code.
Commit Collected from  (Paul et al., 2021) Number of historical commits for the code file in its code change. Excessive history commits may lead developers to fatigue or overconfidence, resulting in their ignorance of security considerations in simpler scenarios.
Complexity Collected from  (Paul et al., 2021; McIntosh et al., 2016) McCabe’s Cyclomatic Complexity (McCabe, 1976). Computed using the Lizard (liz, 2024) analyzer. Code comprehensibility may become more challenging for LLMs as its cyclomatic complexity increases (Paul et al., 2021).
@@
AuthorExperience Collected from  (Paul et al., 2021) Number of closed code commits the author has submitted prior to the commit this code file corresponding to. Experienced authors may write more rigorous and standardized code, while the security vulnerabilities in the code are also relatively more subtle and difficult to detect.
@@
AnnotationHasCode Transformed from  (Chen et al., 2023) Whether the annotations in the code file include code. Code in annotations was abandoned by developers but often interpreted as actual code by LLMs (Chen et al., 2023), leading to incorrect security defect detection.

### v4

@@ 2.2 Data Collection
2.3 Research Procedure of RQ1
Table 1: Versions and hyper-parameters of 6 LLMs studied.
Model Version Ctx Len Parameter
@@ 2.3 Research Procedure of RQ1
2.3.1 LLM Selection
To fully evaluate the performance of LLMs in security defect detection, we selected four representative LLMs from the proprietary GPT and Gemini families, along with two from the open-source Llama 2 family, which are the most popularly used LLMs when we started our experiments in December 2023. Specifically, for proprietary LLMs, we chose GPT-3.5 and GPT-4 provided by the ChatGPT platform (May 2024 version), and GPT-4 Turbo and Gemini Pro accessed through the API. For open-source LLMs, we deployed the smallest and largest chat versions of models in the Llama 2 series–namely, Llama 2 7B and Llama 2 70B–on a server with eight NVIDIA GeForce RTX 4090 GPUs, each with 24 GB of VRAM. To realistically reflect on the practical capabilities of LLMs in security defect detection, we used the default hyperparameter settings throughout our experiments. Some studies mitigate the non-determinism of LLMs by setting the temperature to zero (Atil et al., 2024; Blackwell et al., 2024). In contrast, we refrained from doing so, as this may adversely affect the performance of the LLMs (Song et al., 2024) and may not guarantee determinism (Ouyang et al., 2025; Astekin et al., 2024). We retained the default temperature settings and quantified the non-determinism of LLMs as one aspect of performance evaluation in RQ1.4, rather than artificially suppressing such non-determinism via hyperparameter tuning. The versions, context lengths (Ctx Len), and parameter settings of LLMs used in our study are summarized in Table 1.
@@ 2.3.1 LLM Selection
2.3.2 Prompt Design
Prompt engineering is a process of designing prompts to optimize model performance on downstream tasks (Liu et al., 2023). To design prompts, we followed the best practices published by OpenAI (OpenAI, 2023) and the three prompt design strategies in Zhang et al. (2023): basic prompt, enhanced prompt with auxiliary information, and CoT prompt. During our preliminary exploration stage, we tried prompts with various auxiliary information and different CoT intermediate reasoning steps. To fully harness the capabilities of LLMs in detecting security defects, we finally selected the five best-performing prompts. Fig. 2 illustrates the construction strategies for the five prompts.
@@ 2.3.2 Prompt Design
  • Prompt 4 ($\textbf{P}_{cid}$): $\textbf{P}_{b}$ + Specific CWE Instruction. $\textbf{P}_{cid}$ provides a specific list of the first level CWEs in VIEW-1000 (Research Concepts) [^1] VIEW-1000 is a weaknesses classification framework designed for academic research, widely adopted in prior work (e.g.,  (Pan et al., 2023)). It organizes CWEs into a hierarchical structure based on parent-child relationships among CWE entries. The first level of CWEs in VIEW-1000 encompasses a broad range of commonly encountered security defects, thus applied in Prompt 4 ($\textbf{P}_{cid}$). We acknowledge that providing a list of CWEs may limit the scope of detection set of CWEs and miss security defects that do not fall into the provided CWE categories. However, including all CWEs is infeasible due to the limited context window of the LLM and the potential loss of focus.
@@
Figure 2: Construction templates for the five prompts
Table 2: Adjusted IHMU-category for LLM responses.
Classification Definition Instance
Instrumental (I) The type of the security defect identified in the response is exactly the same as the target security defect, the location is strictly within the code line range of the target security defect, and the defect description can provide effective information for understanding the scenario. Reviewer: (Line 76)I think this is race-y. Let’s say we have a three coroutines scheduled to call function. First one gets to L74 and function yields. Second coroutine gets in before threshold and gets to L68, sets is_deferred and sleeps on L70. First coroutine is finished and unsets on L76. Third coroutine enters still before threshold so L62 doesn’t influence it as planned and gets to L70, sleeps similar time as second coroutine. Then both, second and third coroutines call function in very short time difference.
LLM: **Potential Time of Check to Time of Use (TOCTOU) race condition in ‘Throttler.call()‘** (Lines 58-77):
- The is_deferred flag is used to determine if a function should be deferred. However, multiple threads can check the is_deferred flag almost simultaneously and hence more than one thread might execute the contained code.
- **Solution**: A synchronization mechanism (like a lock) should be employed to ensure thread safety.
Explanation: LLM indicated the security defect identified by the reviewer, with the correct type, accurate code line range and detailed description, thus categorized as Instrumental.
Helpful (H) The type or location of the identified security defect differ to some extent from the target security defect, or the description may be too broad or partially inaccurate. However, the response still provides useful evidence pointing toward the existence of the target defect. Reviewer: (Line 141) If connect_info ever includes a secret of a type other than token (e.g. password) that mask_password looks for, you’d be replacing that with the masked token instead of masking the password or whatever it is. But I guess that’s unlikely.
LLM:**Potential Information Disclosure**(Line 141):
- Logging connection information can potentially expose sensitive details. Even though ‘mask_password‘ is being used, it’s essential to validate that no other sensitive information is leaked.
- **Solution**: Limit the details being logged and ensure that all sensitive data is properly obfuscated before logging.
Explanation: The reviewer indicated that the mask_password function might incorrectly replace different types of sensitive information. While the LLM correctly identified the relevant, it only provided general emphasis on checks for potential sensitive information exposure, failing to recognize this specific risk scenario. Consequently, this instance was categorized as Helpful rather than Instrumental.
Misleading (M) The response does not provide any helpful information, e.g., claiming no defects were found or failing to perform the detection task. Reviewer: (Line 329) This does implicit format detection on the imported disk, which is a well-known security hole. In this specific case a malicious user could create a ’raw’ image which actually contained a malicious qcow2 header specifying a backing file of the the compute host’s root disk. This would result in copying the compute host’s root disk to the specified rbd volume and exposing it to the user.
LLM: Based on the provided code snippet, no security defects are detected. For a more comprehensive analysis, a full code audit in the context of the entire application and its environment would be recommended.
Explanation: LLM did not detect any security defects, but the reviewer actually identified a security hole in the disk format detection. Therefore, this instance was considered Misleading.
@@
Uncertain (U) The response does not mention any defects identified by the reviewer but instead points out the existence of other security defects. Due to the lack of expertise of the researchers about the code and its context, the actual existence of these security defects cannot be confirmed. Reviewer: (Line 469) So not even timeout cancels the future? Will this cause that threads will hang around forever leaking OS resources? This feels dangerous.
LLM: **Potential Insecure Direct Object References (Lines 89-563):**
- The ‘RequestContext‘ class and its methods provide direct access to project and user identifiers, and various methods in this file allow direct manipulation of database and message queue connections. If access controls are not properly enforced or if input validation is weak, this could lead to unauthorized access or manipulation of data.
Explanation: The reviewer identified that the thread handling may lead to OS resource leaks. However, LLM did not mention any about resource leak, instead highlighting a potential improper access issue. Since access mechanisms might be uniformly managed by a middleware, and considering the unavailability of detailed implementation information, we cannot determine whether the issue pointed out by LLM exists. Thus this instance was categorized as Uncertain.
2.3.3 Evaluation Metrics
@@
During the experiments, the responses of LLMs were provided in the form of natural language text, making it necessary to manually inspect these responses to ensure the accuracy of the performance measurements. We considered that LLM-generated responses could be evaluated in a manner similar to the approach adopted by Mahajan et al. (2020) for measuring responses in Stack Overflow. We adapted the evaluation metrics formulated in their work (Mahajan et al., 2020) as follows: First, we excluded the ‘Unavailable’ category defined by Mahajan et al. (2020) because LLMs do not have a case of no response. Next, the definitions of each category were adjusted to match the task of security code review. Specifically, each response was categorized into one of the four rating categories (Instrumental, Helpful, Misleading, Uncertain) detailed in Table 2. To ensure clearer differentiation, for the ‘Helpful’ and ‘Instrumental’ categories, we specified their definitions based on their performance across three key attributes: defect type, location, and description. The corresponding evaluation metrics are I-Score ($\frac{I}{I+H+U+M}\times 100\%$), IH-Score ($\frac{I+H}{I+H+U+M}\times 100\%$), and M-Score ($\frac{M}{I+H+U+M}\times 100\%$). Higher I-Score and IH-Score indicate a stronger ability of the model to detect security defects identified by reviewers. At the same time, a lower M-Score signifies fewer misleading results produced during security defect detection. Given that LLMs exhibit considerable non-determinism, we repeated the experiment for each LLM-prompt combination three times and calculated the average of the three performance scores as the final result, so as to alleviate this threat and ensure a reliable evaluation of LLMs’ capabilities in security defect detection.
2.3.4 Baselines
@@
We tried different tools and learning techniques for detecting security defects to select appropriate baselines for our work.To compare with LLMs, we focused on techniques providing fine-grained detection, including classification, localization, and description of defects. This led us to exclude techniques with insufficient detection granularity, such as ProRLearn (Ren et al., 2024), GRACE (Lu et al., 2024) and LineVul (Fu and Tantithamthavorn, 2022). Given that we evaluated the performance of LLMs on the C/C++ and Python datasets separately, for language-specific static analysis tools, we selected CppCheck (Marjamäki et al., 2024) for C/C++ and Bandit (PyCQA, 2024) for Python. Furthermore, we also used CodeQL (GitHub, 2021) and SonarQube (SA, 2024) to analyze Python files. These two tools can directly analyze code files written in dynamically-typed languages like Python but require code files in statically-typed languages like C/C++ to be compiled first. Therefore, we could not apply them to our C/C++ dataset, which is comprised of uncompiled and isolated C/C++ code files. Additionally, since Semgrep (Dennison et al., 2024) supports multiple programming languages and can analyze individual code files, we adopted it for both C/C++ and Python datasets. To obtain comprehensive analysis results, CodeQL adopted the Python-security-and-quality.qls test suite, covering a wide range of issues from basic code structure and naming conventions to advanced security and performance vulnerabilities. Other tools utilized their default rule sets for analysis.
2.3.5 Data Labelling
@@
Prompts were constructed using five templates and then fed into six LLMs to collect their responses. Since the responses of LLMs were provided in the form of natural language text, to rate the responses accurately, we manually inspected the content of LLM-generated responses, their corresponding source code, and security defects identified by reviewers to categorize them into the four categories formulated in Table 2. To mitigate bias, pilot data labelling was conducted, in which we randomly selected 100 out of 534 code files from our dataset, constructed five prompts for each file and fed these prompts into LLMs to collect responses. Using the definitions of categories in Table 2 as the labelling criteria, the first and third authors labelled the collected responses separately. Discrepancies in labelling results were discussed with the second author to reach a consensus. The inter-rater reliability was calculated using Cohen’s Kappa coefficient (Cohen, 1960), yielding a value of 0.89. Then, the first and third authors labelled all the rest of the responses and discussed any ambiguous cases with the second author to finalize the categorization. Based on the labelling results, we calculated the performance scores of each LLM-prompt combination in the Python and C/C++ datasets. We found that the combination of GPT-4 and prompt $\textbf{P}_{cid}$ outperformed the others (as detailed in Section 3.1). Therefore, we aggregated the responses generated by GPT-4 with $\textbf{P}_{cid}$ across the Python and C/C++ datasets, resulting in a total of 421 responses (some code files exceeded the token limitation of GPT-4) for each of the three repetitive experiments, to further analyze RQ2 and RQ3.
2.3.6 Consistency Calculation
@@
To measure the non-determinism of LLMs, we adopted the metrics in Chang et al. (2024) and utilized entropy to measure the consistency of responses generated by LLMs across three repetitive experiments. Specifically, an LLM generated three responses for a given code file under a specific prompt template. We then calculated the entropy for the four response categories in the three responses. We determined the overall consistency by averaging the entropy across all code files using the following formula: $\frac{1}{R}\sum_{r\in R}\text{entropy}(p^{r}_{1},\dots,p^{r}_{4})$, where R is the response count in each round of the experiment. A higher average entropy suggests the LLM has higher inconsistency in repetitions of experiments.
2.4 Research Procedure of RQ2
@@
We manually inspected the responses of GPT-4 with $\textbf{P}_{cid}$ across the three repetitive experiments to analyze the quality problems present in these responses. We employed the open coding and constant comparison (Glaser, 1965) with a predefined category by the MAXQDA tool for qualitatively analyzing the responses.
@@
Based on the problem categories formulated by Kabir et al. (2024) and the problems of LLM-generated text in various domains identified by previous studies (Borji, 2023; Mitrović et al., 2023; Kalla and Kuraku, 2023), two of the authors firstly inspected a few GPT-4 responses independently and recorded their observations. Then they collaboratively reviewed the inspected responses and discussed to establish an initial codebook, resulting in four themes: Correctness, Understandability, Conciseness, and Compliance, each with a series of problem types. In particular, we excluded the Consistency theme defined in Kabir et al. (2024) from our study. The reason is that the consistency between the LLM’s responses and the reviewers’ assessments (i.e., the ground truth) has been investigated in RQ1. Following the codebook, we conducted a pilot data analysis across two iterations to improve the agreement between the two authors. In each round, ten responses were randomly selected for the two authors to analyze their quality problems. The results were then compared and discussed, with partial adjustments made to the codebook. Since a single response could contain multiple problems, two authors completed the remaining data analysis task where the labels are not mutually exclusive. Therefore, we calculated the agreement level using Krippendorff’s alpha (Krippendorff, 2004) with Jaccard distance, which was improved from 0.40 in the first round to 0.80 in the second, indicating a high level of agreement. Then the two authors completed the remaining data analysis and verified by the second author. Throughout the entire procedure, we employed a negotiated agreement approach (Campbell et al., 2013): any conflicts were consulted and addressed by three authors, ensuring the reliability of data analysis. For each type of problem, we computed its average frequency in three repetitive experiments as the final occurrence rate of this problem in GPT-4 responses.
We consider the finalized codebook to cover all types of quality problems present in the LLM-generated responses within our dataset. To further analyze the proportion of hallucinations in LLMs’ responses, we adopted the granular taxonomy proposed by Huang et al. (2025), which categorizes hallucinations into two main types: Factual and Faithfulness. Factual Hallucination refers to outputs that contradict real-world facts or lack factual grounding, including Factual Contradiction and Factual Fabrication. Faithfulness hallucination, on the other hand, refers to deviations of the output from the instruction, current context, and reasoning process, including Instruction Inconsistency, Context Inconsistency, and Logical Inconsistency. We mapped the quality problems defined in our codebook to these hallucination subtypes to calculate their distribution across the dataset. Detailed mapping and results are provided in Section  3.2.
2.5 Research Procedure of RQ3
@@
To answer RQ3, we adopted the approach formulated by Harrell (Harrell, 2001) to construct a regression model. From Table 2, we can see that the response variables in our dataset are the ratings of I, H, U, and M. In terms of the helpfulness of the response, by definition, I is more important than H, and so forth. Accounting for the ordinal nature of the response variable and the non-determinism of the LLM in three rounds of experiments, we ultimately aggregated the data from all three experiments to fit a cumulative link mixed model (Powers and Xie, 2008). This model supports ordered categorical response variables and allows us to include random effects to group responses in each round together. The specific construction process is described in the following subsections.
@@
Table 3: Factors that may influence the performance of LLM in security code review
@@
Token Adjusted from  Paul et al. (2021); McIntosh et al. (2016) Number of tokens of the code file. Computed using the TikToken OpenAI (2024) tokenizer. Larger code files may be more difficult for LLMs to analyze.
@@
SecurityDefectType Transformed from  Chen et al. (2023) Type of the security defect that the reviewers identified in this code file, i.e., one of the 15 types of security defect formulated by Yu et al. Yu et al. (2023). ChatGPT is proven to display distinct detection performances across various vulnerabilities Chen et al. (2023); Zhou et al. (2023). Thus, we assume that LLMs may be more adept at detecting certain specific types of security defects in code.
Commit Collected from  Paul et al. (2021) Number of historical commits for the code file in its code change. Excessive history commits may lead developers to fatigue or overconfidence, resulting in their ignorance of security considerations in simpler scenarios.
Complexity Collected from  Paul et al. (2021); McIntosh et al. (2016) McCabe’s Cyclomatic Complexity McCabe (1976). Computed using the Lizard Terry (2024) analyzer. Code comprehensibility may become more challenging for LLMs as its cyclomatic complexity increases Paul et al. (2021).
@@
AuthorExperience Collected from  Paul et al. (2021) Number of closed code commits the author has submitted prior to the commit this code file corresponding to. Experienced authors may write more rigorous and standardized code, while the security vulnerabilities in the code are also relatively more subtle and difficult to detect.
@@
AnnotationHasCode Transformed from  Chen et al. (2023) Whether the annotations in the code file include code. Code in annotations was abandoned by developers but often interpreted as actual code by LLMs Chen et al. (2023), leading to incorrect security defect detection.